# Deep Reinforcement Learning Enhanced Rate-Splitting Multiple Access for Interference Mitigation

### v1

@@ IV-D Interference Channel Upper Bounds
In [16] and [17], the authors suggest upper bounds for single antenna interference channels. For the single antenna case, we will use the upper bound given in [17, (58)-(71)] for weak and mixed interference conditions. In this context, for $i=1,2$, let $\bm{H}_{i}=h_{i}$ and $\bm{G}_{j}=g_{j}$. Then, the signal-to-noise ratio (SNR) of user $i$, $\mathrm{SNR}_{i}$, and interference-to-noise ratio (INR) for ${UE}_{i}$ caused by $UE_{j}$, $\mathrm{INR}_{i}$, of [17, (58)-(71)] respectively become $\mathrm{SNR}_{i}=\left|h_{i}\right|^{2}P_{i}/N_{0}$ and $\mathrm{INR}_{i}=\left|g_{j}\right|^{2}P_{j}/N_{0}$. Furthermore, the weak interference condition is defined as $\mathrm{INR}_{1}<\mathrm{SNR}_{2}$ and $\mathrm{INR}_{2}<\mathrm{SNR}_{1}$, while the mixed interference condition encompasses scenarios where either $\mathrm{INR}_{1}\geq\mathrm{SNR}_{2}$ and $\mathrm{INR}_{2}<\mathrm{SNR}_{1}$ or $\mathrm{INR}_{1}<\mathrm{SNR}_{2}$ and $\mathrm{INR}_{2}\geq\mathrm{SNR}_{1}$. Finally, the strong interference condition occurs, when $\mathrm{INR}_{1}>\mathrm{SNR}_{2}$ and $\mathrm{INR}_{2}>\mathrm{SNR}_{1}$. The interference channel becomes equivalent to a compound multiple access channel, in which both users decode all messages. In this case the symmetric rate point is given in [17, (33)] and the whole region is expressed in [16, (5)].
With these definitions, for each channel realization, we first determine whether the channel is a weak, mixed or strong interference channel and then directly apply the corresponding bounding equations in [17, (58)-(71)] or [16, (5)] for the SISO case.
@@ V Simulation Results
The instantiation of the MADDPG algorithm was conducted within the PyTorch 1.9.1 framework. The training architecture incorporates four fully connected layers for both the critic and the actors, providing a robust and sophisticated foundation for the learning process. To assess the algorithm’s convergence, extensive monitoring was conducted over 12,000 episodes for MIMO, 4,000 episodes for MISO, and 2,400 episodes for SISO cases, with each episode consisting of 200 time steps. This comprehensive evaluation resulted in the acquisition of rates across varying SNR scenarios. Within the simulated environment, the channel coefficients $\bm{H}_{i}$, and $\bm{G}_{i}$ for $i=1,2$ are assumed to follow an independent and identically distributed circularly symmetric complex Gaussian distribution with zero mean and unit variance. Irrespective of the antenna configuration, be it SISO, MISO, or MIMO, the total power $P_{i}$ in (16c) is consistently set to 1.
@@
Figure 7: Evolution of the weighted sum-rate for MADDPG with rate-splitting for $M=3$ and $Q=1$ when $SNR=10$ dB and for varying $\beta$ defined in (16a). The skip from one dot to the next represents 100 episodes of training, with the dots appearing after a delay of 500 episodes.
In Fig. 7, we illustrate the evolution of the weighted sum rate under a 10 dB SNR for $M=3$ and $Q=1$ for different weight parameter values; i.e. $\beta$ defined in (16a). In the figure, the skip from one dot to the next represents 100 episodes of training, with dots appearing after a delay of 500 episodes. The training process involves averaging our model over 100 runs, each comprising 1000 time steps realizations. The green outer box, corresponds to the no interference case, which is the ultimate upper bound for the achievable rate region. We also plot the upper bound [18] rate pair for $\beta=0.5$ with a star marker. The learning curve evolution analysis for different $\beta$ values allows us to explore how the rate region is obtained during the learning process. By observing how the system’s performance changes with different $\beta$ values, we gain insights into the algorithm’s behavior across a range of rate configurations, aiding in the assessment of its adaptability and responsiveness to changes in user rate priorities.
@@
In Fig. 9, we systematically investigate the impact of a fixed value of imperfection over SNR, as defined in (23). The estimation error is set as $\frac{(10^{-0.6})}{5}\mathcal{CN}(0,\sigma^{2})$ with $\sigma^{2}=1$. In this analysis, the robustness of the proposed RSMA framework is highlighted. MADDPG with RSMA continues to follow the upper bound even under channel estimation errors. This quality becomes particularly pronounced when compared against alternative schemes, including ZF, MRT, leakage-based precoding, and non-rate-splitting MADDPG. In scenarios with imperfect channel state information at the transmitter (CSIT), the conventional schemes—ZF, MRT, leakage-based precoding, and non-rate-splitting MADDPG— encounter challenges in maintaining their performance due to their limited adaptability to variations in channel estimation errors. As channel estimation error variance is fixed with respect to SNR, the gap between perfect channel state knowledge and imperfect CSIT curves increase with increasing SNR for leakage based precoding, MADDPG with no rate-splitting and ZF. MRT relies on a straightforward strategy of amplifying the signals by the ratio of the received signal strength to the noise level, regardless of the specifics of the channel state. This approach minimizes the dependence on precise channel knowledge, making MRT less susceptible to the challenges associated with obtaining and utilizing accurate real-time channel state information. However, the simplicity that renders MRT resilient also contributes to its suboptimal nature. By adhering to a basic amplification strategy, MRT may not exploit the full potential of channel knowledge for optimizing transmission performance. However, the unique rate-splitting mechanism in MADDPG exhibits exceptional proficiency in alleviating the deleterious effects of channel estimation errors and continues to perform very close to the upper bound.
Additionally, Fig. 10 explores the scenario where imperfection varies according to $\frac{(SNR^{-0.6})}{5}\mathcal{CN}(0,\sigma^{2})$, with the same $\sigma^{2}=1$. Although a fixed channel estimation error is a more limiting assumption, this scenario is more practical. If SNR is high, it is very likely that it is high both for the channel estimation phase and for the data transmission phase. We observe that MADDPG with rate-splitting, ZF and MRT all converge with their imperfect CSIT versions for high SNR. MADDPG with no rate-splitting and leakage based precoder also converge but at very high SNR. Both Figs. 9 and 10 confirm the inherent resilience presented by RSMA, showcasing the superiority of the proposed framework under diverse SNR and channel estimation error conditions.
@@ VI Conclusion and Future Work
Future work includes extending deep reinforcement learning models to handle scenarios with multiple users and designing new algorithms that take energy efficiency into account. Also, exploring hierarchical reinforcement learning approaches to tackle the complexity of rate-splitting optimization, and investigating meta-learning techniques to reduce the need for extensive training in new scenarios are of interest. In short, the future of using deep reinforcement learning for rate-splitting in wireless communications holds great potential for optimizing resource allocation, improving spectral and/or energy efficiency, and enabling more dynamic and adaptable wireless networks. Our work is a first step towards addressing these challenges and will lead to significant advancements in wireless communication systems.

### v2

@@ IV-D Interference Channel Upper Bounds
In [18] and [19], the authors suggest upper bounds for single antenna interference channels. For the single antenna case, we will use the upper bound given in [19, (58)-(71)] for weak and mixed interference conditions. In this context, for $i=1,2$, let $\bm{H}_{i}=h_{i}$ and $\bm{G}_{j}=g_{j}$. Then, the signal-to-noise ratio (SNR) of user $i$, $\mathrm{SNR}_{i}$, and interference-to-noise ratio (INR) for ${UE}_{i}$ caused by $UE_{j}$, $\mathrm{INR}_{i}$, of [19, (58)-(71)] respectively become $\mathrm{SNR}_{i}=\left|h_{i}\right|^{2}P_{i}/N_{0}$ and $\mathrm{INR}_{i}=\left|g_{j}\right|^{2}P_{j}/N_{0}$. Furthermore, the weak interference condition is defined as $\mathrm{INR}_{1}<\mathrm{SNR}_{2}$ and $\mathrm{INR}_{2}<\mathrm{SNR}_{1}$, while the mixed interference condition encompasses scenarios where either $\mathrm{INR}_{1}\geq\mathrm{SNR}_{2}$ and $\mathrm{INR}_{2}<\mathrm{SNR}_{1}$ or $\mathrm{INR}_{1}<\mathrm{SNR}_{2}$ and $\mathrm{INR}_{2}\geq\mathrm{SNR}_{1}$. Finally, the strong interference condition occurs, when $\mathrm{INR}_{1}>\mathrm{SNR}_{2}$ and $\mathrm{INR}_{2}>\mathrm{SNR}_{1}$. The interference channel becomes equivalent to a compound multiple access channel, in which both users decode all messages. In this case the symmetric rate point is given in [19, (33)] and the whole region is expressed in [18, (5)].
With these definitions, for each channel realization, we first determine whether the channel is a weak, mixed or strong interference channel and then directly apply the corresponding bounding equations in [19, (58)-(71)] or [18, (5)] for the SISO case.
@@ V Simulation Results
The instantiation of the MADDPG algorithm was conducted within the PyTorch 1.9.1 framework. The training architecture incorporates four fully connected layers for both the critic and the actors, providing a robust and sophisticated foundation for the learning process. To assess the algorithm’s convergence, extensive monitoring was conducted over 12,000 episodes for MIMO, 4,000 episodes for MISO, and 2,400 episodes for SISO cases, with each episode consisting of 200 time steps. This comprehensive evaluation resulted in the acquisition of rates across varying SNR scenarios. Within the simulated environment, the channel coefficients $\bm{H}_{i}$, and $\bm{G}_{i}$ for $i=1,2$ are assumed to follow an independent and identically distributed circularly symmetric complex Gaussian distribution with zero mean and unit variance. Irrespective of the antenna configuration, be it SISO, MISO, or MIMO, the total power $P_{i}$ in (16b) is consistently set to 1.
@@
Figure 6: The evolution of the rate pair $(R_{1},R_{2})$ during training episodes for MADDPG with rate-splitting when $M=3$, $Q=1$, $SNR=10$ dB and different values of $\beta$ as defined in (16a). Each dot represents 100 episodes of training, with the first dot appearing after 500 episodes. Subsequent dots are plotted every 100 episodes.
In Fig. 6, we illustrate the evolution the rate pair $(R_{1},R_{2})$ as the number of episodes increases. This is examined under 10 dB SNR for $M=3$ and $Q=1$ for different weight parameter values; i.e. $\beta$ defined in (16a). In the figure, the skip from one dot to the next represents 100 episodes of training, with dots appearing after a delay of 500 episodes. The training process involves averaging our model over 100 runs, each comprising 1000 time steps realizations. The green outer box corresponds to the trivial outer bound with no interference, where each user can attain its individual capacity when there is no interference. We also include the rate pair point marked with a star to illustrate a point on the upper bound of the capacity region as defined by [20]. The learning curve evolution analysis for different $\beta$ values allows us to explore how the rate region is obtained during the learning process. By observing how the system’s performance changes with different $\beta$ values, we gain insights into the algorithm’s behavior across a range of rate configurations, aiding in the assessment of its adaptability and responsiveness to changes in user rate priorities.
@@
In Fig. 8, we systematically investigate the impact of a fixed value of imperfection over SNR, as defined in (23). The estimation error is set as $\frac{10^{-0.6}}{5}\mathcal{CN}(0,\sigma^{2})$ with $\sigma^{2}=1$. In this analysis, the robustness of the proposed RSMA framework is highlighted. MADDPG with RSMA continues to follow the upper bound even under channel estimation errors. This quality becomes particularly pronounced when compared against alternative schemes, including ZF, MRT, leakage-based precoding, and non-rate-splitting MADDPG. In scenarios with imperfect channel state information at the transmitter (CSIT), the conventional schemes—ZF, MRT, leakage-based precoding, and non-rate-splitting MADDPG— encounter challenges in maintaining their performance due to their limited adaptability to variations in channel estimation errors. As channel estimation error variance is fixed with respect to SNR, the gap between perfect channel state knowledge and imperfect CSIT curves increase with increasing SNR for leakage based precoding, MADDPG with no rate-splitting and ZF. MRT relies on a straightforward strategy of amplifying the signals by the ratio of the received signal strength to the noise level, regardless of the specifics of the channel state. This approach minimizes the dependence on precise channel knowledge, making MRT less susceptible to the challenges associated with obtaining and utilizing accurate real-time channel state information. However, the simplicity that renders MRT resilient also contributes to its suboptimal nature. By adhering to a basic amplification strategy, MRT may not exploit the full potential of channel knowledge for optimizing transmission performance. However, the unique rate-splitting mechanism in MADDPG exhibits exceptional proficiency in alleviating the deleterious effects of channel estimation errors and continues to perform very close to the upper bound.
Additionally, Fig. 9 explores the scenario where estimation error varies according to $\frac{SNR^{-0.6}}{5}\mathcal{CN}(0,\sigma^{2})$, with $\sigma^{2}=1$. Although a fixed channel estimation error is a more limiting assumption, this scenario is more practical. If SNR is high, it is very likely that it is high both for the channel estimation phase and for the data transmission phase. We observe that MADDPG with rate-splitting, ZF and MRT all converge with their imperfect CSIT versions for high SNR. MADDPG with no rate-splitting and leakage based precoder also converge but at very high SNR. Both Figs. 8 and 9 confirm the inherent resilience presented by RSMA, showcasing the superiority of the proposed framework under diverse SNR and channel estimation error conditions.
@@ VI Conclusion and Future Work
The proposed MADDPG framework could be easily extended to handle scenarios with multiple users and to design new algorithms that take energy efficiency into account. Also, exploring hierarchical reinforcement learning approaches to tackle the complexity of rate-splitting optimization, and investigating meta-learning techniques to reduce the need for extensive training in new scenarios are of interest. In short, the future of using deep reinforcement learning for rate-splitting in wireless communications holds great potential for optimizing resource allocation, improving spectral and/or energy efficiency, and enabling more dynamic and adaptable wireless networks. Our work is a first step towards addressing these challenges and will lead to significant advancements in wireless communication systems.